OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL ^^ Find the area of the shaded region, if the rectangle has dimensions 6 m and 15 m and the circle is inscribed in the rectangle. You may leave your answers in terms of π.

OpenStudy (anonymous):

shaded region of circle or the rectangle without the circle? circle = 3^2 π = 9π m rectangle without the circle = 6 * 15 - 9π = 9(10 - π) m

OpenStudy (anonymous):

rectangle without the circle @KennethNg so that's 90-9π correct? Guess I was right ^^ wanted to verify thanks!~ will fan and give a medal.

OpenStudy (anonymous):

thanks ^^ this is the first time for me to answer a question haha

OpenStudy (anonymous):

cool xD wee~ @KennethNg